# On stabilizers in finite permutation groups

### v2

@@ 2.2. Imprimitive groups, colorings
We first notice that $\mathrm{A\Gamma L}_{1}(8)$ has precisely $4$ orbits on the power set whose stabilizers have orbits of length at most $4$. Let $B=\mathrm{A\Gamma L}_{1}(8)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on a set $\Omega$ of cardinality $32$. It is not hard to see that $B$ has a unique orbit $\Delta^{B}$ on the power set whose setwise stabilizer has orbits of length at most $5$ (actually $4$). Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$. If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting setwise stabilizer will have an orbit of length $4\cdot 2=8$. Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a setwise stabilizer (in $B$) with all orbits of length at most $5$. So there is no hope of getting such a stabilizer in $G$.
@@ 2.3. A result about $3$-colorings
With the same ideas used in the proof of Theorem 2.7, it is possible to obtain the following result concerning $3$-colorings. We observe that Corollary 2.11 answers another question of Babai [Bab22, Question 7c].
@@
Let $G$ be a solvable permutation group on the finite set $\Omega$. Then there exist at least $5$ orbits of $3$-colorings of $\Omega$ whose stabilizers have all orbits of length at most $2$.
@@
We have to prove that $\ell_{2,3}(G)\geq 5$. We can suppose that $G$ is transitive. Primitive groups can be handled with Lemma 2.6. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|=a\geq 2$. Let $G\leqslant B\wr_{a}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis, we have that $\ell_{2,3}(B)\geq 5$. Arguing as in the proof of Theorem 2.7, using these binary $3$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the regular $5$-colorings of $\Omega_{1}$, we can construct $\ell_{1,5}(A)$ binary $3$-colorings of $\Omega$. Finally, from Lemmas 2.4 and 2.5, we have $\ell_{1,5}(A)\geq 5$. ∎
@@
As already noticed by Babai, not even a generic $2$-group can be obtained with just two colors: take $\mathop{\mathrm{Sym}}(4)\wr K$ for any solvable $K$ that is not a $2$-group.
2.4. A CFSG-free proof for arbitrary primitive groups
The foundation stone to prove a result like Theorem 1.1 is the fact that a primitive solvable group of large degree has a regular orbit on the power set [Glu83]. With the trivial exceptions of the alternating and symmetric groups, the same is actually true in the nonsolvable case, but the known proofs of Cameron, Neumann, Saxl and Seress [CNS84, Ser97] require the classification of the finite simple groups. We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] has the following consequence.
@@ 2.4. A CFSG-free proof for arbitrary primitive groups
Let $G$ be a primitive group not containing $\mathop{\mathrm{Alt}}(n)$. With only finitely many exceptions, $G$ has a regular orbit on the power set.
@@
Let $n$ be the degree of $G$. We mimic the short proof in [CNS84], which uses the notions of base size and minimal degree in a clever way. At the crucial point, they obtain $|G|\geq 2^{\sqrt{n/2}}$, and use the classification to remain only with Johnson and Hamming groups, which are handled separetely. While Babai’s theorem [Bab81] is not enough, [SW16, Corollary 1.6] is sufficient for this job, having a thresold of $2^{n^{1/3}(\log n)^{7/3}}$. ∎
Let $G$ be an imprimitive group. To glue together the various pieces in Theorem 2.12, one has to be careful that large alternating groups are not induced on any subset. This fact makes the possible statements more complicated. One consequence is the following:
@@
Let $G$ be a finite permutation group on $\Omega$ which does not involve $\mathop{\mathrm{Alt}}(n)$ for any $n\geq n_{0}$. Then there is a subset of $\Omega$ whose setwise stabilizer has all orbits of length bounded by a function of $n_{0}$.
@@ 3. Primitive groups with solvable stabilizer
We would like to remark that, the idea of fixing a bounded number of points to obtain something interesting, is completely hopeless in imprimitive groups. This is a key difference with respect to setwise stabilizers (see Corollaries 1.2 and 2.11, for example).
@@
Let $B$ be a transitive group with stabilizer $B_{0}$, and let $C_{k}$ be the cyclic group on $k$ points. Let $G=B\wr_{k}C_{k}$ act imprimitively on $|B:B_{0}|k$ points. The stabilizer of a point in $G$ is $H\cong B_{0}\times B^{k-1}$. For each $b\geq 1$, the intersection of any $b$ conjugates of $H$ will contain a copy of $B^{k-b}$, showing that fixing less than $k$ points is useless.
Ultimately, the goal of this section is to prove Theorem 1.4. In the case of primitive solvable groups, this is [LS24, Theorem 1.1]. The proof of this part is really about finding a good centralizer in every irreducible solvable linear group, which was done by Yang [Yan09] using ideas in [MW04]. In general, as explained by Li and Zhang in [LZ11, Theorem 1.1], a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
@@ 3.1. Almost simple groups
Theorem 1.3 will follow from the tables in [LZ11, Sec. 10], together with a few properties of finite simple groups and linear groups. We report a strong version of the Schreier conjecture, which says that the group $\mathop{\mathrm{Out}}(T)$ of the outer automorphisms of a finite simple group is solvable.
@@
This follows from the classification of the finite simple groups. In particular, see [CCN+85, Pag. xvi, Table 5]. For finite simple groups of Lie Type, $\mathop{\mathrm{Out}}(T)$ can be described as an extension of three groups: diagonal, field and graph automorphisms (the field and graph automorphisms commute). These groups are abelian except for the graph automorphisms group of $\mathrm{P}\Omega^{+}_{8}(q)$, which is isomorphic to $\mathop{\mathrm{Sym}}(3)$. ∎
@@
Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag. 4]. We have that $H$ is an extension of $H_{0}=H\cap G_{0}$ with
@@
To deal with (iii) uniformly, we use the fact that a solvable linear group has derived length bounded by a function on the rank [Dix68]. The (relatively high) constant $10$ is attained by the Fischer group $Fi_{{}_{23}}$, which makes smooth the remaining computations. We point out some mistakes in [LZ11, Table 15], and rely instead on the [GAP4] library CTblLib (v1.3.9). [^2] One mistake is in correspondence of the solvable maximal subgroups of $Fi_{{}_{23}}$. For example, a potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer. With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}.(S4\times 2S4)$. A group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case. ∎
A weaker version of Theorem 1.3 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple. It would be interesting to find a proof of this fact which does not rely on the classification of the finite simple groups.
@@ 3.2. Groups of product type
$T\wr_{d}S\>\leqslant\>G\>\leqslant\>X\wr_{d}S.$
(See also [Bur21, Sec. 8.2].) Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action. Let $W\cong Y^{d}$ be the subgroup of $V$ such that $H=WS\cap G$. Stabilizing two points corresponds to intersecting $H$ with one of its conjugates. To prove Theorem 1.4, we will find $v\in G\cap V$ such that the derived length of
@@
is bounded. The proof will follow, because $H\cap H^{v}\leqslant I_{v}$, and the derived length of a subgroup is bounded by the derived length of the group.
Proof of Theorem 1.4.
Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$. Since $Ker(\pi)=G\cap V$, we have
$dl(I_{v})\leq dl(I_{v}\cap V)+dl(\pi(I_{v})).$
@@
Now $dl(Y)$ is bounded by Theorem 1.3, and the rest of the proof deals with $\pi(I_{v})$. It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be quite bigger. For every $w\in W$, $\sigma\in S$ and $v\in V$, an explicit computation gives
@@
we obtain
(3.1) $\pi(I_{v})=\{\sigma\in S:\>x_{\sigma(i)}\in Yx_{i}Y\mbox{ for each }i=1,\ldots,d\}.$
We are going to look at the double cosets of $Y$ as colors. Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets. For any $v\in V$, and $j\in\{1,\ldots,r\}$, let
@@
(3.2) $\pi(I_{v})=\cap_{j=1}^{r}Stab_{S}(\Delta_{j})=\cap_{j=1}^{r-1}Stab_{S}(\Delta_{j}).$
To construct $v$, we will use only two elements $c_{1}$ and $c_{2}$ (and so two sets $\Delta_{1}$ and $\Delta_{2}=\{1,\ldots,d\}\setminus\Delta_{1}$). One reason is that we want $v$ to lie in $G\cap V$, and not just in $V$. This is assured if we choose representatives $c_{i}$’s which are in $T$. We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus\{1\}$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is solvable. Finally, choosing $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ as provided by Corollary 1.2, from (3.2) we get $dl(\pi(I_{v}))\leq 3$, concluding $dl(I_{v})\leq 13$ as desired. ∎
@@
In the proof of Theorem 1.4, it might seem wasteful to construct $v$ with only two elements $c_{1}$ and $c_{2}$. As explained, one reason is that we want $v\in G$, but there is also a different motivation. In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist many doubly transitive groups of this type (as $\mathop{\mathrm{Sym}}(5)$, for example). A more detailed analysis of highly transitive almost simple groups, and the additional use of Corollary 2.11, might help to find the right constant in Theorem 1.4. We remark that this constant is not even known in the case of solvable groups, where it lies somewhere between $4$ and $9$ (see [LS24, Remark 2.9]).
@@
Let $\{x_{1},x_{2}\}\subseteq\Omega$ be as provided by Theorem 1.4. If $\{x_{1},x_{2}\}\subseteq\Delta$, then $C_{G}(\Delta)$ is solvable of bounded derived length. If $|\Delta|\leq 4$, then $Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $Stab_{G}(\Delta)$ is solvable of bounded derived lenght. ∎

### v3

@@ 2.3. Imprimitive groups
We first notice that $\mathrm{A\Gamma L}_{1}(8)$ has precisely $4$ orbits on the power set whose stabilizers have orbits of length at most $4$, and $4$ is attained by one of these orbits (a crucial fact is that $\ell_{2,3}(\mathrm{A\Gamma L}_{1}(8))=3$ in Table 1). Let $B=\mathrm{A\Gamma L}_{1}(8)\wr_{4}\mathop{\mathrm{Sym}}(4)$ act imprimitively on a set $\Omega$ of cardinality $32$. It is not hard to see that $B$ has a unique orbit $\Delta^{B}$ on the power set where the set-stabilizer has orbits of length at most $5$ (actually $4$). Now let $G=B\wr_{2}\mathop{\mathrm{Sym}}(2)$ act on $\Omega\times\{\alpha,\beta\}$. If we color both $\Omega\times\{\alpha\}$ and $\Omega\times\{\beta\}$ using $\Delta$, then the resulting set-stabilizer will have an orbit of length $4\cdot 2=8$. Otherwise, we are forced to color a subset, say of $\Omega\times\{\alpha\}$, which does not provide a set-stabilizer (in $B$) with all orbits of length at most $5$. So there is no hope of getting such a stabilizer in $G$.
@@ 2.4. A result about $3$-colorings
With the same ideas developed in the proof of Theorem 2.10, we obtain the following powerful result concerning $3$-colorings.
@@
Let $G$ be a solvable permutation group on the finite set $\Omega$, $|\Omega|\geq 2$. Then there exist at least $5$ orbits of $3$-colorings of $\Omega$ whose stabilizers have all orbits of length at most $2$.
@@
We have to prove that $\ell_{3,2}(G)\geq 5$. We work by induction on $|\Omega|$, and we can suppose that $G$ is transitive. Also, primitive groups can be handled with Lemma 2.7. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis, we have that $\ell_{3,2}(B)\geq 5$. We argue as in Subsection 2.3, with $x=5$ and $y=3$. Using the $2$-asymmetric $3$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the asymmetric $5$-colorings of $\Omega_{1}$, we can construct $\ell_{5,1}(A)$ inequivalent $2$-asymmetric $3$-colorings of $\Omega$. We have $\ell_{5,1}(A)\geq 5$ from Corollary 2.6, which makes the induction smooth and completes the proof. ∎
The correspondent structural result answers [Bab22, Question 7c].
@@
See [Glu24] for another positive result (with two colors) that emphatizes the importance of the prime $2$. In general, five colors are necessary and sufficient to achieve a trivial stabilizer [Ser96, Theorem 1.2].
2.5. CFSG-free results for nonsolvable groups
The foundation stone to prove a result like Theorem 1.1 is the fact that a primitive solvable group of large degree has a regular orbit on the power set [Glu83]. With the trivial exceptions of the alternating and symmetric groups, the same is actually true in the nonsolvable case, but the known proofs of Cameron, Neumann, Saxl and Seress [CNS84, Ser97] require the Classification of Finite Simple Groups. Actually, Cameron [Cam86] proved a stronger statement concerning the number of regular orbits. We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] is strong enough to provide an easier proof of Cameron’s theorem. [^1] Babai pointed out that a slightly weaker observation, which gives an elementary proof of [CNS84], was made by himself in [Bab22, Theorem 15.1].
@@ 2.5. CFSG-free results for nonsolvable groups
Let $G$ be a primitive group of degree $n$ not containing $\mathop{\mathrm{Alt}}(n)$. Then the probability that a random orbit on the power set is a regular orbit tends to $1$ as $n$ grows.
@@
With the notation in [Cam86], it is sufficient to prove that $\varepsilon_{1}(G)\to 0$ when $n$ grows. We mimic Cameron’s proof, which uses the notions of base size and minimal degree in a clever way. At the crucial point, he obtains the desired result provided that $|G|\leq c^{\sqrt{n}}$, where $c$ is a constant (see the Remark at the end of page 308). Then, the Classification of Finite Simple groups is used to bound the order of a primitive group, remaining only with Johnson and Hamming groups, which are handled separately. While Babai’s theorem [Bab81] is not enough, [SW16, Corollary 1.6] is sufficient for this job, having a thresold of $d^{\,n^{1/3}(\log n)^{7/3}}$ where $d$ is a constant. ∎
We stress that Babai’s theorem [Bab81] has several applications, and Theorem 2.15 is a nice example where Sun-Wilmes improvement turns out to be crucial. (See [Ebe24] for another recent example.) We are now ready to prove Theorem 1.3, which answers [Bab22, Question 6a]. Again, we prove a slightly stronger statement.
@@
There exists a constant $C$ such that the following holds. Let $G$ be a permutation group on the finite set $\Omega$, $|\Omega|\geq 2$. Then there are at least two orbits of subsets of $\Omega$ whose setwise stabilizers have all nonalternating composition factors of order at most $C$.
Proof.
After Theorem 2.15, it is easy to see that such a constant $C$ exists for all primitive groups. Given $C$, we call good the orbits on the power set with the property in the statement. We can suppose that $G$ is transitive. By the Jordan-Hölder theorem, the property of having all nonalternating composition factors of order at most $C$, is a group property preserved by group extensions. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By induction, $B$ admits two good orbits on $\mathcal{P}(\Omega_{2})$ (equivalently, two good $2$-colorings). Arguing as in Subsection 2.3 (set $x=y=2$), using these good orbits of $B$ as colors for $\Omega_{1}$, and the two good $2$-colorings of $\Omega_{1}$, we can construct two inequivalent good orbits of $G$ on $\mathcal{P}(\Omega)$. The proof is complete. ∎
With more work, and using the Classification of Finite Simple Groups, it should be possible to find an explicit version of Theorem 2.16 (it is enough to find $C$ for all primitive groups).
@@ 3. Primitive groups with solvable stabilizer
In this section, we move to consider pointwise stabilizers of very small sets. We would like to remark that, the idea of fixing a bounded number of points to obtain something interesting, is completely hopeless in imprimitive groups. This is a key difference with respect to stabilizing sets (see Corollaries 1.2 or 2.14, for example).
@@
Let $B$ be a transitive group with stabilizer $B_{0}$, and let $S\leqslant\mathop{\mathrm{Sym}}(d)$. Let $G=B\wr_{d}S$ act imprimitively on $|B:B_{0}|d$ points. The stabilizer of a point in $G$, say $H$, contains a copy of $B_{0}\times B^{d-1}$. For each $b\geq 1$, the intersection of any $b$ conjugates of $H$ will contain a copy of $B^{d-b}$, showing that fixing less than $d$ points is useless.
Therefore, we will focus on primitive groups, with the goal of proving Theorem 1.5. This proof is more algebraic, but the combinatorial ideas in the previous section will also play a key role. In the case of primitive solvable groups, Theorem 1.5 is [LS24, Theorem 1.1]. This part is really about finding a good centralizer in every irreducible solvable linear group, which was done by Yang [Yan09] using ideas in [MW04]. In general, as explained by Li and Zhang in [LZ11, Sec. 2], it is not hard to see that a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
@@ 3.1. Almost simple groups
Theorem 1.4 will follow from the tables in [LZ11, Sec. 10], together with a few properties of finite simple groups and linear groups. We first report a strong version of the Schreier conjecture, which says that the group $\mathop{\mathrm{Out}}(T)$ of the outer automorphisms of a finite simple group is solvable.
@@
This follows from the Classification of Finite Simple Groups. In particular, see [CCN+85, Pag. xvi, Table 5]. For finite simple groups of Lie Type, $\mathop{\mathrm{Out}}(T)$ can be described as an extension of three groups: diagonal, field and graph automorphisms (the field and graph automorphisms commute). These groups are abelian except for the graph automorphisms group of $\mathrm{P}\Omega^{+}_{8}(q)$, which is isomorphic to $\mathop{\mathrm{Sym}}(3)$. ∎
@@
Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag. 4]. We have that the maximal subgroup $H$ is an extension of $H_{0}=H\cap G_{0}$ with
@@
To deal with (iii) uniformly, we use the fact that a solvable linear group has derived length bounded by a function on the rank [Dix68]. This and Lemma 3.2 prove Theorem 1.4 for some absolute constant. The (relatively high) constant $10$ is attained by the Fischer group $Fi_{{}_{23}}$, which makes smooth the remaining computations in (i)-(iii). A more careful inspection of the tables shows that $dl(H_{0})\leq 6$ in the infinite families of (iii), while we use a computer for the remaining single groups. We point out some mistakes in [LZ11, Table 15], and rely instead on the [GAP4] library CTblLib (v1.3.9). [^2] One mistake is in correspondence of $Fi_{{}_{23}}$. In fact, $Fi_{{}_{23}}$ has a unique conjugacy class of solvable maximal subgroups, namely AtlasSubgroup(‘‘Fi23’’,7)), with structure $43^{1+8}.2^{1+6}.3^{1+2}.2S$ 4 (with the ATLAS notation) and derived length $10$. A potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer. With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}.(S4\times 2S4)$. Any group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case. ∎
A weaker version of Theorem 1.4 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple. It would be interesting to find a proof of this curious fact which does not rely on the Classification of Finite Simple Groups.
@@ 3.2. Groups of product type
$T\wr_{d}S\>\leqslant\>G\>\leqslant\>X\wr_{d}S,$
and acts on $|X:Y|^{d}$ points. (See also [Bur21, Sec. 8.2].)
Proof of Theorem 1.5.
Let $V\cong X^{d}$ be the base subgroup of $X\wr_{d}S$, and let $H<G$ be the stabilizer of a point in the product action. Let $W\cong Y^{d}$ be the subgroup of $V$ such that $H=WS\cap G$. Stabilizing two points corresponds to intersecting $H$ with one of its conjugates. To prove the theorem, we will find $v\in G\cap V$ such that the derived length of
@@
is bounded by an absolute constant (actually $13$). The proof will follow, because $H\cap H^{v}\leqslant I_{v}$, and the derived length of a subgroup is bounded by the derived length of the group. Let $\pi\colon G\to S$ be the natural projection, and let $v\in V$. Since $Ker(\pi)=G\cap V$, we have
(3.1) $dl(I_{v})\>\leq\>dl(I_{v}\cap V)+dl(\pi(I_{v})).$
@@
Now $dl(Y)\leq 10$ by Theorem 1.4, and the rest of the proof deals with $\pi(I_{v})$. It is not hard to see that $C_{S}(v)=S\cap S^{v}\subseteq\pi(I_{v})$, but in general $\pi(I_{v})$ can be quite bigger. For every $w\in W$, $\sigma\in S$ and $v\in V$, an explicit computation gives
@@
we obtain the key equality
(3.2) $\pi(I_{(x_{1},\ldots,x_{d})})\>=\>\{\sigma\in S:\>x_{\sigma(i)}\in Yx_{i}Y\mbox{ for each }i=1,\ldots,d\}.$
This suggests to look at the double cosets of $Y$ in $X$ as colors, and connects Theorem 1.5 with the matter of Section 2. Let $c_{1},\ldots,c_{r}\in X$ be representatives of these double cosets ($r\geq 2$). For any $v\in V$, and $j\in\{1,\ldots,r\}$, let
@@
(3.3) $\pi(I_{v})\>=\>\cap_{j=1}^{r}Stab_{S}(\Delta_{j})\>=\>\cap_{j=1}^{r-1}Stab_{S}(\Delta_{j}).$
To construct $v$, we will use only two elements $c_{1}$ and $c_{2}$ (and so two sets $\Delta_{1}$ and $\Delta_{2}=\{1,\ldots,d\}\setminus\Delta_{1}$). One reason is that we want $v$ to lie in $G\cap V$, and not just in $V$. This is assured if we choose representatives $c_{i}$’s which are in $T$. We can surely choose $c_{1}=1$ and $c_{2}\in T\setminus\{1\}$, because otherwise $T\leqslant Y$, which is impossible because $Y$ is core-free. Finally, if $\Delta_{1}=\Delta\subseteq\{1,\ldots,d\}$ is as provided by Corollary 1.2, from (3.3) we get $dl(\pi(I_{v}))\leq 3$, concluding by (3.1) that $dl(I_{v})\leq 13$ as desired. ∎
@@
In the proof of Theorem 1.5, it might seem wasteful to construct $v$ with only two elements $c_{1}$ and $c_{2}$. As explained, one reason is that we want $v\in G$, but there is also a different motivation. In fact, it is well known that the number $r$ of double cosets of $Y$ in $X$ equals the rank of the primitive group $X$ with solvable stabilizer $Y$, and there actually exist some doubly transitive groups of this type (as $\mathop{\mathrm{Sym}}(5)$, for example). A more detailed analysis of almost simple groups, and the additional use of Corollary 2.14, might help to find the right constant in Theorem 1.5. In the case of primitive solvable groups, this lies somewhere between $4$ and $9$ (see [LS24, Remark 2.9]). H.Y. Huang pointed out that $5$ is attained by the almost simple group $\mathop{\mathrm{Aut}}(\mathrm{P}\Omega^{+}_{8}(3))$. With some more work, it is not hard to use this information to construct a group of product type where the constant $6$ is achieved.
A small corollary on set-stabilizers can be deduced from Theorem 1.5.
@@
Let $\{x_{1},x_{2}\}\subseteq\Omega$ be as provided by Theorem 1.5. If $\{x_{1},x_{2}\}\subseteq\Delta$, then $C_{G}(\Delta)$ is solvable of bounded derived length. Now $Stab_{G}(\Delta)/C_{G}(\Delta)\leqslant\mathop{\mathrm{Sym}}(4)$, and it follows that $Stab_{G}(\Delta)$ is solvable of bounded derived lenght. ∎